To determine the number of deer in a game presurve, conservationists catch 697 deer and tags them. Later 408 deer are caught; 204 of them are tagged. How many deer in the presurve?
Answer by edjones(8007)

204/408=1/2 the deer are tagged.
697 deer are 1/2 the deer in the preserve.
There are 697*2=1394 deer in the preserve.
.
Ed